OpenStudy (anonymous):

Will give medal for any help :) Question: A pendulum is 18 feet long. Its central angle is 44°. The pendulum makes one back and forth swing every 12 seconds. To the nearest foot, how far does the pendulum swing each minute?

OpenStudy (anonymous):



OpenStudy (anonymous):

I still don't understand. So the opposite side is the distance, right? But how would you find that when you're only given one side?

OpenStudy (anonymous):

the sin of 44=0.69=x/18 so x=18*.69=12.5 this is 1/2 the swing so full swing =25 feet in 12 seconds so how many in a minute? 60/12 =5 5*25=125 feet

OpenStudy (anonymous):

I get it now, thank you so much!